OpenStudy (anonymous):

Keep all answers in fractional form. Solve the systems of equations. y=2x-5, y=6x+7

OpenStudy (anonymous):

You have 2 equations where y = something. Since they both equal y, that also means they equal each other: i.e. \(2x-5 = 6x+7\), so you can just solve for x using algebra. Then you can find y by plugging x into one of the original equations.

OpenStudy (amistre64):

systems of equations is a bit of a step up from the other stuff you were doing ...

OpenStudy (amistre64):

the most basic solution would be to solve for one variable and use that in the other equation

OpenStudy (anonymous):

Yeah... i think this homework is just going to keep on getting harder. But how do i solve for one variable if there is two variables in the equation?

OpenStudy (amistre64):

y=2x-5 ; well, its already solved for y so I spose that part was redundant, so lets substitute this value in the other equation where it has a y y=6x+7 2x-5 = 6x+7 ; now all we have to do is solve for x

OpenStudy (anonymous):

try to express y in terms of x. or x in terms of y. then substitute.

OpenStudy (anonymous):

would the answer be -3?

OpenStudy (anonymous):

That is the value of x, yes! Now find the value of y by plugging the x into one of the original equations.

OpenStudy (amistre64):

dunno lets chk: 2x-5 = 6x+7 ; lets +5 to each side +5 +5 ------------- 2x+0 = 6x+12 ; now we can subtract 6x from each side -6x -6x --------------- -4x+0 = 0+12 ; and lets re write it to clean it up -4x = 12 ; the only thing left to do is divide out the -4 /-4 /-4 -------- x = -3 ; that is the solution for x, now that we have a solid answer we can use it in our original equations to find y

OpenStudy (anonymous):

uh, would it be -11? because y=6x+7 y=6(-3)+7 y=-18+7 y=-11..right?

OpenStudy (amistre64):

since x=-3 y=2x-5 = 2(-3) -5 = -6 -5 = -11 OR y =6x+7 =6(-3)+7 = -18 + 7 = -11 from the looks of it; when x=-3; y=-11

OpenStudy (anonymous):

Yess! I did it right! :D

OpenStudy (amistre64):

great job :)